# An Evaluation of Explanation Methods for Black-Box Detectors of Machine-Generated Text

### v1

@@ 3.3 Evaluation of Usefulness
We define usefulness as the explanation method’s ability to improve users’ understanding of the detector’s behavior. To quantify this, we modify the design for a forward simulation experiment in Hase and Bansal (2020). In phase 1 of the experiment, users inspect decisions of the detector on a set of documents $\{a_{j}\}$ without explanations. In phase 2 they are then instructed to anticipate the detector’s decision (not to guess the true document class) on a second set $\{b_{j}\}$. Phase 3 provides explanations for set $\{a_{j}\}$ (Figure 2). Users conclude the experiment by labeling the documents from set $\{b_{j}\}$ again in phase 4. The change in user accuracy from phase 2 to 4 is reported as a measure of performance. [^4] See Figure 4 in Appendix B for an overview of all phases. We ask three additional questions from Hoffman et al. (2019) for every explanation shown in phase 3 to measure perceived usefulness. These are provided in Appendix C. In order to be able to track whether users can pick up regularities between the different phases, we apply a document selection strategy that ensures a minimal overlap of features in the documents, see Appendix B.
@@ 4.1 Dataset
A subset of the H3 dataset by Guo et al. (2023) is used for the experiments. They add ChatGPT written answers to question answering datasets from various domains. [^5] ELI5 Fan et al. (2019), WikiQA Yang et al. (2015), FiQA Maia et al. (2018), Medical Dialog He et al. (2020) and Wikipedia computer science articles Guo et al. (2023). Published under the CC-BY-SA licence. We remove all questions from WikiQA and ELI5, as the source datasets contain crawling artifacts that make identifying human texts trivial. Moreover, only documents that are between 50 and 150 words long are kept. This is to ensure a sufficient input length for the zero-shot detector. In the end, 1016 documents remain. Due to the high computation time involved with generating explanations, only 30% of documents (N=305) are used (stratified split of machine-generated and human-written documents) for our experiments. As discussed in Section 3, we derive additional synthetic datasets from this set.
4.2 Detectors and Explanation Methods
We apply the explanation methods to three detectors of machine-generated text: The fine-tuned RoBERTa model of Solaiman et al. (2019), the model of Guo et al. (2023) (which uses the same base model, but has been trained on the full H3 dataset), and the zero-shot model of Mitchell et al. (2023), see Appendix A. We perform no further fine-tuning.
@@ 4.2 Detectors and Explanation Methods
Feature Importance Explanation Methods assign scores to individual tokens in the document with the goal to quantify the effect a token has on the decision of the classifier. LIME trains a local surrogate model on a set of data point perturbations and corresponding detector outputs Ribeiro et al. (2016). The number of perturbations to use for fitting this model, and its size, have to be chosen manually. The number of samples was set to 1k for the RoBERTa-based detectors and 500 for the zero-shot method to match SHAP’s runtime. [^6] Resulting in an average of $\approx$30s for the RoBERTa-based detectors and $\approx$415s for DetectGPT per explanation. The default number of 10 features to show as an explanation appears appropriate for the document length used here (50-150 tokens). The implementation of SHAP we use (Partition Explainer: Lundberg and Lee, 2017, default parameters) computes Owen values Owen (1977) as a measure of feature importance.
@@
Rule-based explanation methods – like Anchor, explain complex decision processes with short rules Ribeiro et al. (2018). Among valid rules as outlined in Section 3, Anchor’s search algorithm attempts to select those that apply to the highest proportion of perturbations in the local neighborhood. In its default implementation, Anchor is often unable to terminate for a single document within an hour of runtime with these detectors. We employ the following strategies to make the computation of Anchors feasible here: We choose a low target level of precision ($\tau=0.75$) and impose a limit on the number of samples used during construction (200 samples per candidate Anchor). For generating perturbations, we employ DistillRoBERTa (distilroberta-base 82.8M params; Sanh et al., 2020) instead of DistillBERT, to increase their coherence and edit at most 20% of tokens per perturbation. Combined, this reduces the computation time to an average of roughly 5 minutes per explanation for the RoBERTa-based detectors and 15 minutes for DetectGPT on the original documents.
@@
Perturbation Strategy An experiment that tests the effect of different perturbation strategies (Appendix D) did not single out a method that works equally well for all detectors. In the interest of consistency, perturbations are generated by replacing tokens with the mask token of the detector’s tokenizer for both LIME and SHAP throughout all experiments. Anchor offers masking with a specified token or to perturb with a language model. The latter strategy was used, as it was found to terminate considerably faster with these detectors.
@@ 4.3 User Study
We recruited 36 participants (B.Sc., M.Sc., and PhD students) with a background in computer science and English reading proficiency at the C1 CEFR level or higher. Of them, 27 stated that they had never worked with explanation methods before. The instructions shown to users on the individual explanation methods are provided in Appendix C. We offered a compensation of €10 for their participation. The experiment was conducted online through a purpose-built web service.
@@
Table 1: Results for the faithfulness and stability experiments. The best scores are highlighted in bold.
5 Results
Aggregate results per explanation method for faithfulness and stability are reported in Table 1, those from the user study in Table 2. See Appendix D for results per detector-explanation method pairing.
@@ 5.1 Faithfulness
For the token removal experiment, results for initially correct predictions ($f(d_{i})=y_{i}$) and initially wrong predictions ($f(d_{i})\neq y_{i}$) are plotted separately in Figure 3 to allow for consistent interpretation of accuracy scores Arras et al. (2016). We report the change in accuracy at $k=10$ tokens removed ($\Delta_{k=10}$) in Table 1, corresponding to the maximum number of tokens Anchor and LIME include in their explanations. We provide the mean scores for five explanations with random feature importance vectors as a baseline. For $f(d_{i})=y_{i}$, the most important features are removed first. A faithful (feature importance) explanation method should have a steep drop in accuracy. For $k<10$ tokens removed, the accuracy of LIME drops slightly faster than that of SHAP. The average accuracy drops below 50% for SHAP at roughly 10 tokens masked. For Anchor, the accuracy drops slower than for all other methods.
Figure 3: Accuracy at $k$ tokens masked. A faithful explanation method should feature a steep decline (initially correct predictions, left) or steep incline (initially wrong predictions, right). Only SHAP explanations cover more than 10 tokens. Mean across all detectors, error bars at ±$\pm 1$ standard error.
@@ 5.2 Stability
Consistency For Krippendorff’s $\alpha$, a score of 0 reflects an agreement by chance, 1 perfect agreement, and negative values systematic disagreement Krippendorff (1970). Scoring 0.136 and 0.160 respectively, LIME’s and Anchor’s consistency in re-runs is far below what could be considered reliable. SHAP (Partition Explainer) is deterministic for the detectors of Guo et al. (2023) and Solaiman et al. (2019). Note that the zero-shot detector of Mitchell et al. (2023) is not deterministic. We include this detector in the aggregate scores in Table 1 for consistency with the other experiments. The ranking remains unaffected by this, per-detector results are provided in Appendix D.
Continuity SHAP’s explanations remain stable under small perturbations ($\alpha>0.8$ for the detectors of Guo et al., 2023 and Solaiman et al., 2019). Those from LIME and Anchor do not, but agree with each other better than explanations from the random baseline.
Contrastivity SHAP ranks highest in both synthetic data checks $c_{\text{inter}}$ and $c_{\text{intra}}$: It can identify that the filled-in part is responsible for flipping the label more reliably than LIME and Anchor.
@@ 5.3 Usefulness
Forward Simulation Table 2 shows the change in user accuracy from phase 2 to phase 4 and the results from McNemar’s tests. SHAP is the best-performing method, followed by Anchor and LIME. The increase in user accuracy for both SHAP and Anchor is not significant at $p<0.05$. However, participants who have seen LIME explanations perform 13.12% worse after being shown explanations than they did before (p=0.006).
Perceived Usefulness Conversely, users rated LIME best, SHAP second, and Anchor third across all three questions assessing perceived usefulness. For SHAP, users only tended to agree with the first item, where they were asked whether they could understand why the detector decided the way it did from an explanation (Figure 5 in Appendix C). They disagreed with the two other items, asking whether they better understood how the detector works and whether they thought the information from the explanation would help them perform better in the second round of annotation. They tended to agree for LIME with all three statements, for Anchor to disagree on average.
@@ 6 Discussion
Faithfulness The explanations from SHAP and Anchor are in line with the data generation strategy in the pointing game, and LIME achieves the lowest pointing game accuracies. The pointing game and the token removal experiment are based on different assumptions: The pointing game rewards if the explanation method is consistent with the assumption that a classifier relies on material that is associated with a predicted class more than with unrelated material. This is a reasonable assumption, especially for classifiers with much better than random accuracy. The token removal experiment rewards if the explanation method ranks features on top that if removed in a greedy fashion influence prediction the most. This might be at odds with other notions of importance, such as Shapley values, that estimate importance averaged over different combinations of features.
@@
Stability The partition tree used by SHAP’s Partition Explainer is obtained with a deterministic algorithm. SHAP is therefore able to accomplish a perfect score in the consistency experiment for the two deterministic detectors (in contrast to LIME that uses a different random seed and set of perturbations for each explanation attempt). The difference between SHAP, LIME, and Anchor is less pronounced in the continuity experiment, which measures how explanations are affected by small changes to the input documents. Regarding contrastivity, where we assess whether explanations for similar documents with different predictions are sufficiently different, SHAP is more performant than LIME and Anchor: SHAP attributes importance in the expected way more often than the other methods ($c_{\text{inter}}$). SHAP is also considerably more successful at identifying that the filled-in part, and not the shared part, is responsible for flipping the label ($c_{\text{intra}}$).
@@
Usefulness When measuring the usefulness of explanations presented to users, different aspects must be differentiated: a (partial) understanding of the model to be explained (the detector in our case) vs. the perceived degree of understanding of the model. These two may not be the same, it can be that an explanation model gives plausible explanations that do not provide actual insight into the detector’s behaviour. Another subtle distinction is that between (perceived or actual) understanding of model behaviour vs. understanding of task characteristics (what makes human and generated texts different). We have tested increased model understanding in the forward simulation experiment, where we measure the effect of seeing predictions without and with explanations on the ability to anticipate model behavior on new inputs. We also reported perceived usefulness of the different explanations. However, even though we carefully phrased the questionnaire to inquire the perceived ability to understand model behaviour (I now better understand how the detector works), it is conceivable that users conflated that with a better understanding of task characteristics.
Our results show that none of the methods substantially increased the users’ ability to anticipate detector predictions (with SHAP and Anchor having a small positive but not statistically significant effect). One reason could be that in contrast to simpler tasks for which feature-based explanation methods had been applied to before (e.g., binary sentiment prediction), the possible feature space is much larger for the MGT detection task (obvious sentiment-bearing words vs. subtle frequency choices or constructions). Good performance in the automated metrics ( faithfulness and stability) did not translate to usefulness in the forward simulation. The method with the best perceived usefulness (LIME) shows the worst performance there (as well as in many of the automated metrics). Therefore, our study is a strong motivation to rethink the applicability of local feature-based explanations to complex tasks, and highlights the importance of accompanying automated evaluation of explanation methods with user studies.
@@ 7 Conclusion
In this work, we conduct the first evaluation of explanation methods for detectors of machine-generated text. We find that SHAP fulfills the theoretically motivated properties of good explanations of faithfulness, stability, and usefulness best in our experiments. However, no explanation method led to a significant increase in performance in a user study that tested users’ ability to predict the behavior of detectors with the help of explanations. For LIME, users’ perceived usefulness of explanations did not align with the measured performance, nor with the results from the faithfulness and stability experiments.
Given the results from the user study, we advise against implementing combinations of these methods and detectors in their current form into systems that face untrained users. We do however see their exploratory value for model- or dataset inspection. Based on the results from the faithfulness and stability experiments, we recommend SHAP for this application.
8 Limitations
@@
Appendix A Appendix: Detectors
Guo (Guo et al., 2023) With their dataset, Guo et al. also ship a fine-tuned RoBERTa model (roberta-base 125M params; Liu et al., 2019) . The accuracy of this detector on the sub-set for which explanations were generated is 0.99.
Solaiman (Solaiman et al., 2019) This detector was fine-tuned on the Webtext and GPT-2 output datasets Solaiman et al. (2019), but uses the same base model as Guo et al. (2023). It is therefore treated as an out-of-distribution detector. Its accuracy is 0.92.
DetectGPT (Mitchell et al., 2023) This is a zero-shot method. It is set up here with a smaller language model (pythia-70m: Biderman et al., 2023) as suggested by Mireshghallah et al. (2023) and accomplishes an accuracy of 0.74. This, and reducing the number of perturbations per evaluation from 100 to 5 is done to reduce inference time from roughly 6.3 (15.8 with GPT-2) to 0.9 seconds per document.777 The default setting for this detector would require $10^{5}$ generations with GPT-2 for a single explanation. A comparison with the original implementation is provided in Appendix D.
Appendix B Appendix: User Study
Document Selection

### v2

@@ 3.3 Evaluation of Usefulness
We define usefulness as the explanation method’s ability to improve users’ understanding of the detector’s behavior. Note that we do not aim to measure users’ performance at the base task of detecting machine-generated text. Instead, our goal is to measure their ability to predict the detector’s decision, both with and without explanations. To do so, we modify the design for a forward simulation experiment in Hase and Bansal (2020). As in their design, in phase 1 of the experiment, users inspect decisions of the detector on a set of documents $\{a_{j}\}$ without explanations. In phase 2 they are then instructed to anticipate the detector’s decision (not to guess the true document class) on a second set $\{b_{j}\}$. Phase 3 provides explanations for set $\{a_{j}\}$ (Figure 4). Users conclude the experiment by labeling the documents from set $\{b_{j}\}$ again in phase 4. The change in user accuracy from phase 2 to 4 is reported as a measure of performance. We randomize example ordering between participants.
Figure 4: User Study. Information shown for LIME in phase 3 (left); and in the annotation phases 2 and 4 (right).
While relying on this single performance measure makes it difficult to isolate the direct impact of explanations from users’ growing familiarity with features indicative of human-written or machine-generated text, our setup is designed to reflect typical real-world model inspection: In such contexts, users need to examine numerous explanations to adequately learn about a classifier. Crucially, this design still allows for a fair relative comparison of different explanation methods.
We add three additional questions from Hoffman et al. (2019), asked for every explanation shown in phase 3 to measure perceived usefulness. These are provided in Appendix A. In order to be able to track whether users can pick up regularities between the different phases, we apply a document selection strategy that ensures a minimal overlap of features in the documents, see Appendix A.
@@ 4.1 Dataset
We use a subset of the H3 dataset by Guo et al. (2023) for our experiments. They add ChatGPT written answers to question answering datasets from various domains (ELI5: Fan et al., 2019, WikiQA: Yang et al., 2015, FiQA: Maia et al., 2018, Medical Dialog: He et al., 2020 and Wikipedia computer science articles: Guo et al., 2023; published under the CC-BY-SA license). We remove all questions from WikiQA and ELI5, as the source datasets contain crawling artifacts that make identifying human texts trivial. Moreover, only documents that are between 50 and 150 words long are kept. This is to ensure a sufficient input length for the zero-shot detector. In the end, 1016 documents remain. Due to the high computation time involved with generating explanations, only 30% of documents (N=305) are used (stratified split of machine-generated and human-written documents) for our experiments. As discussed in Section 3, however, we derive additional synthetic datasets from this set, and repeat all experiments with 3 detectors.
4.2 Detectors
We apply the explanation methods to three detectors of machine-generated text:
Guo Guo et al. (2023).
With their dataset, Guo et al. also ship a fine-tuned RoBERTa model (roberta-base 125M: Liu et al., 2019) . The accuracy of this detector on the sub-set for which explanations were generated is 0.99.
Solaiman Solaiman et al. (2019).
This detector was fine-tuned on the Webtext and GPT-2 output datasets Solaiman et al. (2019), but uses the same base model as Guo et al. (2023). It is therefore treated as an out-of-distribution detector. Its accuracy is 0.92.
DetectGPT Mitchell et al. (2023).
This is a zero-shot method. It is set up here with a smaller language model (pythia-70m: Biderman et al., 2023) as suggested by Mireshghallah et al. (2023) and accomplishes an accuracy of 0.74. This, and reducing the number of perturbations per evaluation from 100 to 5 is done to reduce inference time from roughly 6.3 (15.8 with GPT-2) to 0.9 seconds per document. The default setting for this detector would require $10^{5}$ generations with GPT-2 for a single explanation. A comparison with the original implementation is provided in Appendix C.
4.3 Explanation Methods
Feature importance explanation methods
@@ 4.3 Explanation Methods
assign scores to individual tokens in the document with the goal to quantify the effect a token has on the decision of the classifier. LIME trains a local surrogate model on a set of data point perturbations and corresponding detector outputs Ribeiro et al. (2016). The number of perturbations to use for fitting this model, and its size, have to be chosen manually. The number of samples was set to 1k for the RoBERTa-based detectors and 500 for the zero-shot method to match SHAP’s runtime. This results in an average of $\approx$30s for the RoBERTa-based detectors and $\approx$415s for DetectGPT per explanation. The default number of 10 features to show as an explanation (Figure 4) appears appropriate for the document length used here (50-150 tokens).
The implementation of SHAP we use (Partition Explainer; Lundberg and Lee, 2017; default parameters) computes Owen values Owen (1977) as a measure of feature importance.
Rule-based explanation methods
@@
– like Anchor, explain complex decision processes with short rules Ribeiro et al. (2018). Among valid rules as outlined in Section 3, Anchor’s search algorithm attempts to select those that apply to the highest proportion of perturbations in the local neighborhood. In its default implementation, Anchor is often unable to terminate for a single document within an hour of runtime with these detectors. We employ the following strategies to make the computation of Anchors feasible here: We choose a low target level of precision ($\tau=0.75$) and impose a limit on the number of samples used during construction (200 samples per candidate Anchor). For generating perturbations, we employ DistillRoBERTa (distilroberta-base: 82.8M params Sanh et al., 2020) instead of DistillBERT, to increase their coherence and edit at most 20% of tokens per perturbation. Combined, this reduces the computation time to an average of roughly 5 minutes per explanation for the RoBERTa-based detectors and 15 minutes for DetectGPT on the original documents.
Perturbation strategy.
@@
An experiment that tests the effect of different perturbation strategies (Appendix B) did not single out a method that works equally well for all detectors. In the interest of consistency, perturbations are generated by replacing tokens with the mask token of the detector’s tokenizer for both LIME and SHAP throughout all experiments. Anchor offers masking with a specified token or to perturb with a language model. The latter strategy was used, as it was found to terminate considerably faster with these detectors.
@@ 4.4 User Study
We recruited 36 participants (B.Sc., M.Sc., and PhD students) with a background in computer science and English reading proficiency at the C1 CEFR level or higher between 20/03/2024–12/04/2024. Of them, 27 stated that they had never worked with explanation methods before. The instructions shown to users on the individual explanation methods are provided in Appendix A. The participants were offered a compensation of €10 for their participation. The experiment was conducted online through a purpose-built web service, through which we also documented informed consent. We use the default presentation formats of the respective explanation methods to ensure that our analysis is informative about existing applications on MGT detection (e.g., Mosca et al., 2023; Liu et al., 2023; Yu et al., 2023).
5 Results
Table 1: Aggregate results for the faithfulness and stability experiments. Mean scores across the three detectors used for evaluation. The best scores are highlighted in bold.
@@ 5 Results
Aggregate results per explanation method for faithfulness and stability are reported in Table 1, those from the user study in Table 2. See Appendix B for results per detector-explanation method pairing. We provide the mean scores for five explanations consisting of random feature importance vectors as a baseline.
@@ 5.1 Faithfulness
For the token removal experiment, results for initially correct predictions ($f(d_{i})=y_{i}$) and initially wrong predictions ($f(d_{i})\neq y_{i}$) are plotted separately in Figure 5 to allow for consistent interpretation of accuracy scores Arras et al. (2016). We report the change in accuracy at $k=10$ tokens removed ($\Delta_{k=10}$) in Table 1, corresponding to the maximum number of tokens Anchor and LIME include in their explanations. For $f(d_{i})=y_{i}$, the most important features are removed first. A faithful (feature importance) explanation method should show a steep drop in accuracy. For $k<10$ tokens removed, the accuracy of LIME drops slightly faster than that of SHAP. The average accuracy drops below 50% for SHAP at roughly 10 tokens masked. For Anchor, the accuracy drops slower than for all other methods.
Figure 5: Accuracy at $k$ tokens masked. A faithful explanation method should feature a steep decline (initially correct predictions, left) or steep incline (initially wrong predictions, right). Only SHAP explanations cover more than 10 tokens. Mean across all detectors, error bars at $\pm 1$ standard error.
@@ 5.2 Stability
Consistency.
For Krippendorff’s $\alpha$, a score of 0 reflects an agreement by chance, 1 perfect agreement, and negative values systematic disagreement Krippendorff (1970). Scoring 0.136 and 0.160 respectively, LIME’s and Anchor’s consistency in re-runs is far below what could be considered reliable. SHAP (Partition Explainer) is deterministic for the detectors of Guo et al. (2023) and Solaiman et al. (2019). Note that the zero-shot detector of Mitchell et al. (2023) is not deterministic. We include this detector in the aggregate scores in Table 1 for consistency with the other experiments. The ranking remains unaffected by this, per-detector results are provided in Appendix B.
Continuity.
SHAP’s explanations remain stable under small perturbations ($\alpha>0.8$ for the detectors of Guo et al., 2023 and Solaiman et al., 2019). Those from LIME and Anchor do not, but agree with each other better than explanations from the random baseline.
Contrastivity.
SHAP ranks highest in both synthetic data checks $c_{\text{inter}}$ and $c_{\text{intra}}$: It can identify that the filled-in part is responsible for flipping the label more reliably than LIME and Anchor.
@@ 5.3 Usefulness
Forward simulation.
Table 2 shows the change in user accuracy from phase 2 to phase 4 and the results from McNemar’s tests. SHAP is the best-performing method, followed by Anchor and LIME. The increase in user accuracy for both SHAP and Anchor is not significant at $p<0.05$. However, participants who have seen LIME explanations perform 13.12% worse after being shown explanations than they did before (p=0.006).
Perceived usefulness.
Conversely, users rated LIME best, SHAP second, and Anchor third across all three questions assessing perceived usefulness. For SHAP, users only tended to agree with the first item, where they were asked whether they could understand why the detector decided the way it did from an explanation (annotator instructions provided in Appendix A). They disagreed with the two other items, asking whether they better understood how the detector works and whether they thought the information from the explanation would help them perform better in the second round of annotation. They tended to agree for LIME with all three statements, for Anchor to disagree on average.
@@ 6 Discussion
6.1 Faithfulness
The explanations from SHAP and Anchor are in line with the data generation strategy in the pointing game, and LIME achieves the lowest pointing game accuracies. The pointing game and the token removal experiment are based on different assumptions: The pointing game rewards if the explanation method is consistent with the assumption that a classifier relies on material that is associated with a predicted class more than with unrelated material. This is a reasonable expectation, especially for classifiers with much better than random accuracy. The token removal experiment rewards if the explanation method ranks features on top that if removed in a greedy fashion influence prediction the most. This might be at odds with other notions of importance, such as Shapley values, that estimate importance averaged over different combinations of features. Furthermore, it remains unclear whether a steeper initial drop in accuracy, as observed for LIME, or an overall higher drop in accuracy, as observed for SHAP, is more desirable (see left plot in Figure 5), especially given that explanation length is ultimately a design choice: LIME explanations show only the 10 most influential words by default, whereas SHAP provides scores for every word in the input in our setup.
@@ 6.1 Faithfulness
6.2 Stability
The partition tree used by SHAP’s Partition Explainer is obtained with a deterministic algorithm. SHAP is therefore able to accomplish a perfect score in the experiment for consistency (across re-runs) for the two deterministic detectors (in contrast to LIME that uses a different random seed and set of perturbations for each explanation attempt). This is precisely why we included two additional measures of stability that rely on convergence to a lesser extent: in fact, the difference between SHAP, LIME, and Anchor is less pronounced in the continuity experiment, which measures how explanations are affected by small changes to the input documents. Here, the assumption is that similar documents – from the perspective of the detector – should have similar explanations. We use a two step-process to obtain similar perturbations: first, we obtain semantically similar documents by querying language models to provide partial reformulation. In contrast to explanations for vision tasks (e.g., Alvarez Melis and Jaakkola, 2018; where continuous perturbations can be chosen such that they have a limited impact on the classifier output to be explained), perturbation in discrete input domains such as text can have more severe impacts on the prediction. We limit this unwanted effect by discarding perturbations that would flip the label prediction. However, filtering by label flip is just a proxy for similarity from the detector’s perspective, and does not differentiate between different types of effects that could lead to the same predicted label: some reformulations may feature low similarity to the original text (and necessitate dissimilar explanations to retain faithfulness), but be assigned the same label by the detector. Consequently, our continuity experiments may underestimate the agreement of explanations (high agreement equals high continuity).
Regarding contrastivity, where we assess whether explanations for similar documents with different predictions are sufficiently different, SHAP is more performant than LIME and Anchor: SHAP attributes importance in the expected way more often than the other methods ($c_{\text{inter}}$). SHAP is also considerably more successful at identifying that the filled-in part, and not the shared part, is responsible for flipping the label ($c_{\text{intra}}$).
@@ 6.2 Stability
6.3 Usefulness
When measuring the usefulness of explanations presented to users, different aspects must be differentiated: a (partial) understanding of the model (the detector in our case) vs. the perceived degree of understanding of the model. These two may not be the same, it can be that an explanation model gives plausible explanations that do not provide actual insight into the detector’s behaviour. Another subtle distinction is that between (perceived or actual) understanding of model behaviour vs. understanding of task characteristics (what makes human and generated texts different). We have tested for model understanding in the forward simulation experiment, where we measure the effect of seeing predictions without and with explanations on the ability to anticipate model behavior on new inputs. We also reported perceived usefulness of the different explanations. However, even though we carefully phrased the questionnaire to inquire the perceived ability to understand model behaviour (I now better understand how the detector works), it is conceivable that users conflated this with better understanding of task characteristics.
Our results show that none of the methods substantially increased the users’ ability to anticipate detector predictions (with SHAP and Anchor having a small positive but not statistically significant effect). One reason could be that in contrast to simpler tasks for which feature-based explanation methods had been applied to before (e.g., binary sentiment prediction), the possible feature space is much larger for the MGT detection task (obvious sentiment-bearing words vs. subtle frequency choices or constructions). Good performance in the automated metrics ( faithfulness and stability) did not translate to usefulness in the forward simulation. The method with the best perceived usefulness (LIME) shows the worst performance there (as well as in many of the automated metrics). We hypothesize that this might be partially due to differences in explanation length: LIME explanations are limited to the 10 most influential words (see Figure 4) whereas SHAP provides importance scores for every word in the input. Given the between-subject setup of our user study, an analysis of this factor will need to be explored in future work. Our study is a strong motivation to rethink the applicability of local feature-based explanations to complex tasks, and highlights the importance of accompanying automated evaluation of explanation methods with user studies.
@@ 7 Conclusion
In this work, we conduct the first evaluation of explanation methods for detectors of machine-generated text. We find that SHAP fulfills the theoretically motivated properties of good explanations of faithfulness, stability, and usefulness best in our experiments. However, no explanation method led to a significant increase in performance in a user study that tested users’ ability to predict the behavior of detectors with the help of explanations. Notably, for LIME, users’ perceived usefulness of explanations did not align with the measured decrease in user’s performance, nor with the results from the faithfulness and stability experiments. Given the results from the user study, we advise against implementing combinations of these methods and detectors in their current form into systems that face untrained users. We do however see their exploratory value for model- or dataset inspection. Based on the results from the faithfulness and stability experiments, we recommend SHAP for this application.
We have focused on providing a relative comparison between explanation methods, and gathered insights into the suitability of feature importance explanations for tasks with complex feature sets. We did not however, aim to provide insights into specific detectors or to document useful features for MGT detection itself, as that would have required a dedicated analysis of individual explanation method- and detectors pairings, which is beyond the scope of our this work.
Ultimately, our study serves to demonstrate the necessity of evaluating model-agnostic explanation methods before applying them to new tasks and models. While our study can only partially indicate the general suitability of feature importance as an explanation format for language-model-based classifiers, it points to the need for further evaluation: specifically, questions remain about when feature importance provides sufficient insights and for which tasks it is most effective. We advocate for evaluation across additional tasks to better understand the conditions under which feature importance can be an adequate explanation format.
Limitations
@@
We restrict our study to three representative language-model based detectors and one popular benchmark dataset, due to resource constraints imposed by the inclusion of a user study.
Acknowledgments
This research has been funded by the Vienna Science and Technology Fund (WWTF)[10.47379/VRG19008] "Knowledge-infused Deep Learning for Natural Language Processing".
Appendix A User study
Document selection.